The perimeter of a rectangle is 54 inches. The width of the rectangle is 8 inches.
What s the length of the rectangle? Explain your answer.

Answer by macston(5194)

Perimeter of a rectangle=2(L+W) where L is the length and W is the width
54 inches=2(L+8 inches)
54 in=2L+16 in Subtract 16 in from each side
38 in=2L divide each side by 2
19 in=L ANSWER Length is 19 in
CHECK
P=2(L+W)
54 in=2(19 in+8 in)
54 in=2(27 in)
54 in=54 in